OpenStudy (studyhoney):

please check 3 questions

OpenStudy (studyhoney):

11. Factor the GCF: 28a^3b^4 + 20a^2b^2 − 16ab^3. 4ab(7a^2b + 5a − 4b2) 4ab2(7a^2b + 5a − 4b) 4ab2(7a^2b^2 + 5a − 4b) 4ab(7a^2b^3 + 5a − 4b) ^ i think this is C 13. Determine the factors of x^2 − 7x − 10. (x + 2)(x − 5) Prime (x − 2)(x + 5) (x + 10)(x − 1) ^ i think this is B 14. The equation f(x) = 4x^2 − 16x + 9 represents a parabola. What is the vertex of the parabola? (2, −7) (−2, 57) (4, 9) (−4, −137) ^ i think this is A

OpenStudy (reemii):

11. C I think yo'ure right :-)

OpenStudy (reemii):

14 correct

OpenStudy (reemii):

For 13, 'Prime' is "no integer zeros" ?

OpenStudy (studyhoney):

@reemii C is right for 11 just checked it again :)

OpenStudy (mathstudent55):

All are correct.

OpenStudy (reemii):

@studyhoney yes, I corrected my post . good job

OpenStudy (mathstudent55):

Prime means not factorable.

OpenStudy (reemii):

\((x-x_1)(x-x_2)\) with \(x_1,x_2 = \frac{7\pm \sqrt{89}}{2}\) is not an option then?

OpenStudy (mathstudent55):

Factorable means rational factors. Irrational factors make the polynomial considered unfactorable, or prime.